an airplane makes a 990 km flight with a tailwind and returns, flying into the same wind.The total flying time is 3hrs 20 mins and the airplanes speed in still air is600 km/h
what is the speed of the wind.
I have the total distance travelled to be 990+990=1980
Total time 200 min rtn or 100 min each way
Then I am stuck

Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(30993)

an airplane makes a 990 km flight with a tailwind and returns, flying into the same wind.The total flying time is 3hrs 20 mins and the airplanes speed in still air is600 km/h
what is the speed of the wind.
I have the total distance travelled to be 990+990=1980
Total time 200 min rtn or 100 min each way
****************
It's not 100 mins each way. Downwind ground speed is greater than upwind.
-----------
d = distance one way
d/(600 + w) + d/(600 - w) = 10/3 hours
990/(600 + w) + 990/(600 - w) = 10/3 hours
Multiply by 3*(660 + w)*(600 - w)
2970*(600 - w) + 2970*(600 + w) = 10*(600 + w)*(600 - w)
297*1200 = 360000 - w^2
w^2 = 360000 - 356400 = 3600
w = 60 km/hr

Answer by lwsshak3(6522)

an airplane makes a 990 km flight with a tailwind and returns, flying into the same wind.The total flying time is 3hrs 20 mins and the airplanes speed in still air is 600 km/h
what is the speed of the wind.
**
let x=speed of wind
600+x=speed of plane with tail wind
600-x=speed of plane into wind
Total flying time=3hr+20min=20+1/3 hr=10/3 hr
Travel time = distance/speed
..
990/(600+x)+990/(600-x)=10/3
LCD:3(600+x)(600-x)
3*990(600-x)+3*990(600+x)=10(600+x)(600-x)
3*990(600+x+600-x)=10(360000-x^2)
(3*990*1200)/10=360000-x^2
356400=360000-x^2
x^2=360000-356400
x^2=3600
x=√(3600)
x=60
speed of wind=60 km/hr